OpenStudy (anonymous):

Help? Solve the following system of equations. x + y = 12 x – y = 6 A.(3, 9) B.no solution C.infinite number of solutions D.(9, 3)

pooja195 (pooja195):

So we have 2 equations x+y=12 <---equation 1 x-y=6 <----equation 2 I know the substitution method well enough to teach it to you so lets stick with that :)) Lets start by isolating a variable equation 1 has a variable alone lets isolate it x+y=12 <----equation 1 Subtract y from both sides x=-y+12 Now we have x=-y+12 <---equation 1 x-y=6<---equation 2 Now we know the value of x let's plug equation #1 into equation #2 \[\huge~\rm~ -y+12-y=6\]Solve for y tell me what you get :)

OpenStudy (anonymous):

Is it C?

pooja195 (pooja195):

No

OpenStudy (anonymous):

I need help :/ I'm in 8th grade and math is hard.

pooja195 (pooja195):

Im a person who helps people who are willing to learn i guess we are both stuck

OpenStudy (anonymous):

Well can you give me a hint?

pooja195 (pooja195):

−y+12−y=6 Combine like terms -y-y=? basically saying -1-1=?

OpenStudy (anonymous):

No solution??

pooja195 (pooja195):

No -.-

OpenStudy (anonymous):

Don't yell at me!

pooja195 (pooja195):

o.O i yelled at you?

OpenStudy (anonymous):

No he did. Whoever that person is

pooja195 (pooja195):

Ok -1-1=?

OpenStudy (anonymous):

-2?

OpenStudy (alexandervonhumboldt2):

yes -1-1=-2

OpenStudy (anonymous):

Okay!

pooja195 (pooja195):

Good −y+12−y=6 -2y+12=6 now we subtract 12 from both sides 6-12=?

OpenStudy (anonymous):

-6

OpenStudy (alexandervonhumboldt2):

yes

pooja195 (pooja195):

Good -2y=-6 divide both sides by -2 -2/-2=1 so just y -6/-2=?

OpenStudy (anonymous):

-8

pooja195 (pooja195):

We're dividing not adding

pooja195 (pooja195):

-6 divided by -2 =?

OpenStudy (anonymous):

3

pooja195 (pooja195):

Good so y=3 only one of your options has y as 3 what is the answer?

OpenStudy (anonymous):

I would go with A

pooja195 (pooja195):

Nope (x,y) y=3 (x,3)

pooja195 (pooja195):

Try again

OpenStudy (anonymous):

I got it! Thanks!

pooja195 (pooja195):

What did you get?